OpenStudy (anonymous):

Find the derivative of the function:

OpenStudy (anonymous):

you need a function for that

OpenStudy (anonymous):

f(x)= [(4x^9)+(e^(3x+8))] / (5x^9) +2

OpenStudy (anonymous):

explain steps

OpenStudy (anonymous):

omg, why?

OpenStudy (anonymous):

Thats gonna be a mess! :/

OpenStudy (ybarrap):

Possible derivation: d/dx((4 x^9+e^(3 x+8))/(5 x^9)+2) | Differentiate the sum term by term and factor out constants: = | 1/5 (d/dx((4 x^9+e^(3 x+8))/x^9))+d/dx(2) | The derivative of 2 is zero: = | 1/5 (d/dx((4 x^9+e^(3 x+8))/x^9))+0 | Use the product rule, d/dx(u v) = v ( du)/( dx)+u ( dv)/( dx), where u = 1/x^9 and v = 4 x^9+e^(3 x+8): = | 1/5 ((d/dx(4 x^9+e^(3 x+8)))/x^9+(4 x^9+e^(3 x+8)) (d/dx(1/x^9))) | The derivative of 1/x^9 is -9/x^10: = | 1/5 ((d/dx(4 x^9+e^(3 x+8)))/x^9+(4 x^9+e^(3 x+8)) (-9/x^10)) | Differentiate the sum term by term and factor out constants: = | 1/5 ((4 (d/dx(x^9))+d/dx(e^(3 x+8)))/x^9-(9 (4 x^9+e^(3 x+8)))/x^10) | Use the chain rule, d/dx(e^(3 x+8)) = ( de^u)/( du) ( du)/( dx), where u = 3 x+8 and ( de^u)/( du) = e^u: = | 1/5 ((4 (d/dx(x^9))+e^(3 x+8) (d/dx(3 x+8)))/x^9-(9 (4 x^9+e^(3 x+8)))/x^10) | The derivative of x^9 is 9 x^8: = | 1/5 ((e^(3 x+8) (d/dx(3 x+8))+4 (9 x^8))/x^9-(9 (4 x^9+e^(3 x+8)))/x^10) | Differentiate the sum term by term and factor out constants: = | 1/5 ((e^(3 x+8) (3 (d/dx(x))+d/dx(8))+36 x^8)/x^9-(9 (4 x^9+e^(3 x+8)))/x^10) | The derivative of 8 is zero: = | 1/5 ((e^(3 x+8) (3 (d/dx(x))+0)+36 x^8)/x^9-(9 (4 x^9+e^(3 x+8)))/x^10) | The derivative of x is 1: = | 1/5 ((36 x^8+3 1 e^(3 x+8))/x^9-(9 (4 x^9+e^(3 x+8)))/x^10)

OpenStudy (anonymous):

how did you type that so fast?

OpenStudy (anonymous):

Lol he copied it from Wolframalpha

OpenStudy (anonymous):

Scroll down to "derivative" and there is a "show steps" button.

myininaya (myininaya):

\[y=\frac{4x^9+e^{3x+8}}{5x^9+2}\] Take natural log of both sides! \[\ln(y)=\ln(\frac{4x^9+e^{3x+8}}{5x^9+2})\] Using properties of log we can rewrite the left hand side as follows: \[\ln(y)=\ln(4x^9+e^{3x+8})-\ln(5x^9+2)\] Now take derivative of both sides to find y': \[\frac{y'}{y}=\frac{4 \cdot 9 \cdot x^{9-1}+(3x+8)'e^{3x+8}}{4x^9+e^{3x+9}}-\frac{5 \cdot 9 \cdot x^{9-1}+0}{5x^9+2}\] so we have \[\frac{y'}{y}=\frac{36x^{8}+(3+0)'e^{3x+8}}{4x^9+e^{3x+9}}-\frac{45x^{8}}{5x^9+2}\] \[\frac{y'}{y}=\frac{36 x^{8}+3e^{3x+8}}{4x^9+e^{3x+9}}-\frac{45 x^{8}}{5x^9+2}\] To find y' multiply y on both sides: \[y'=y (\frac{36 x^{8}+3e^{3x+8}}{4x^9+e^{3x+9}}-\frac{45 x^{8}}{5x^9+2})\] and we replace y with \[\frac{4x^9+e^{3x+8}}{5x^9+2}\]